caira12:

Erica solved the equation −5x − 25 = 78; her work is shown below. Identify the error and where it was made. (5 points) −5x − 25 = 78 Step 1: −5x − 25 − 25 = 78 − 25 Step 2: −5x = 53 Step 3: = -5x/-5 = 53/-5 Step 4: x = -53/5 Select one: a. Step 1: She should have added 25 to each side. b. Step 1: She should have divided both sides by −5. c. Step 3: She should have divided both sides by −5. d. Step 3: She should have multiplied both sides by −5.

Logic007:

\(-5x-25=78\)

Logic007:

Step 1: Add both sides by 25

Logic007:

\(-5x-25+25=78+25\)

Logic007:

Step 2: Simplify it \(-5x=103\)

Logic007:

Step 3 : Divide both sides by -5

Logic007:

\(\frac{-5}{-5}x=\frac{103}{-5}\)

Logic007:

Step 4: Simplify it \(x=\frac{103}{-5}\)

caira12:

so step 3?

Logic007:

Option A is correct bcoz she should hv added 25 to each side

caira12:

ohhhhh

caira12:

thank you

Logic007:

u can try to compare my answers with the question. np ^